OpenStudy (anonymous):

simplify 5 + 22 - 3(6) ÷ 3

OpenStudy (anonymous):

21

OpenStudy (anonymous):

we have : 5+22-3*(6)/6 first answer to :- 3*6 = -18 . then : -18/3 =- 6 have : 5+22-6=27-6 ;)

OpenStudy (anonymous):

its not 22 its the 2 and the 2 exponent.

OpenStudy (anonymous):

2^2

OpenStudy (anonymous):

yeah that!

OpenStudy (anonymous):

5+2^2-3(6)/3 5+4-18/3 5+4-6 9-6 3

OpenStudy (anonymous):

K ! So first answer to : 2^2= ? ? = 4 then : -3*6=-18 and -18/3=-6 then have : 5+4-6=9-6 =>3 This order is important ;)